OpenStudy (anonymous):

***HELP***MEDAL WILL BE GIVEN*** How do i change x^2+25x+144 into the formula f(x) = a(x - h)^2 + k?

OpenStudy (anonymous):

x^2 + 2*(25/2)x + (25/2)^2 -(25/2)^2 +144 (x+ (25/2))^2 -(49/4)

OpenStudy (anonymous):

@dv4a Wait what.

OpenStudy (anonymous):

you must complete the square. do you know how to do that?

OpenStudy (anonymous):

i used the (a+b)^2 formula

OpenStudy (anonymous):

@dolphinluv21 some what but i just learned it can you elaborate? @dv4a so i plug my a and b in and where does that work?

OpenStudy (anonymous):

you need to make x^2+25x into a (a+b)^2 formula (a+b)^2 = a^2 + 2*a *b + b^2 i simply assumed 25x to be 2 * (25/2) { the 2 cancels out so it's the same thing} so now a = x and b = 25/2

OpenStudy (anonymous):

because you dont have b^2 you simply insert a b^2 ie (25/2)^2 and then subtract it.

OpenStudy (anonymous):

first you take 144 and subtract it from both sides. so your equation right now is (x^2 + 25x + __) = -144 to find the number where the underscores are... you simply take the 25 that was given and divide it by 2 and then square it. That gives you 156.25. so your equation is now x^2 + 25x + 156.25 = -144 + 156.25 (you MUST add the 156.25 to both sides) so now to simplify... (x + 12.5)^2 + 12.25 = f(x)

OpenStudy (anonymous):

@dolphinluv21 so wait x^2+25+156.25= 12.25 you woulndt subtract 12.25

OpenStudy (anonymous):

@dolphinluv21

OpenStudy (anonymous):

*correction: it should have been + 49/4 my bad -.- and as for subtracting 12.25 you are only subtracting so that there is no value in the right hand side and only in the left hand side (if that is what you are asking)

OpenStudy (anonymous):

@dv4a no because he or she got (x + 12.5)^2 + 12.25 = f(x) but thats not subtracting 12.25 from 156.25

OpenStudy (anonymous):

i think i was right before he is wrong in subtracting he should have : x^2 + 25x + 156.25 = -144 + 156.25 x^2 + 25x +156.25 -12.25 = 12.25 -12,25

OpenStudy (anonymous):

*not wrong

OpenStudy (anonymous):

oh yes that is my bad it should be -12.25! Sorry for the confusion.